The roots of the equation 3x2 + 4x – 5 = 0 are α, β. Find the values of the following:

  1. 1/α + 1/β
  2. α2 + β2    
1
Expert's answer
2021-03-16T08:40:17-0400
"3x^2+4x-5=0"

"a=3, b=4, c=-5"

By Vieta theorem


"\alpha+\beta=-\dfrac{b}{a}"

"\alpha\beta=\dfrac{c}{a}"

1)


"\dfrac{1}{\alpha}+\dfrac{1}{\beta}=\dfrac{\alpha+\beta}{\alpha\beta}"

"=\dfrac{-\dfrac{b}{a}}{\dfrac{c}{a}}=-\dfrac{b}{c}=\dfrac{4}{5}"

2)


"\alpha^2+\beta^2=(\alpha+\beta)^2-2\alpha\beta"

"=(-\dfrac{b}{a})^2-2(\dfrac{c}{a})"

"=(-\dfrac{4}{3})^2-2(\dfrac{-5}{3})=\dfrac{46}{9}"